OpenStudy (anonymous):

evaluate (a^4)^3 when a = (2r3)^1r6 Ill rewrite it neater

OpenStudy (anonymous):

\[\huge(a ^{4})^{3} \]

OpenStudy (anonymous):

\[ \huge a= (\frac{ 2 }{ 3 })^{\frac{ 1 }{ 6 }}\]

OpenStudy (anonymous):

evaluate

OpenStudy (anonymous):

i just sub a equal whatever in a^12 but i can never get it right

jimthompson5910 (jim_thompson5910):

(a^4)^3 = a^(4*3) (a^4)^3 = a^(12) So if you want to find (a^4)^3, then you can find a^(12) instead since they are the same

jimthompson5910 (jim_thompson5910):

\[ \Large a= (\frac{ 2 }{ 3 })^{\frac{ 1 }{ 6 }}\] \[ \Large a^{12}\] \[ \Large \left((\frac{ 2 }{ 3 })^{\frac{ 1 }{ 6 }}\right)^{12}\] \[ \Large (\frac{ 2 }{ 3 })^{\frac{ 1 }{ 6 } * 12}\] \[ \Large (\frac{ 2 }{ 3 })^{\frac{ 12 }{ 6 }}\] \[ \Large (\frac{ 2 }{ 3 })^{2}\] I'll let you finish

OpenStudy (anonymous):

so the final answer is 4r9?

OpenStudy (anonymous):

wow thanks so much anyways

jimthompson5910 (jim_thompson5910):

you mean 4/9 right?

jimthompson5910 (jim_thompson5910):

not sure how/where the r comes into play

jimthompson5910 (jim_thompson5910):

btw \[\Large \frac{4}{9} = 4/9\]

OpenStudy (anonymous):

oh lol in my calc to write something over something its something r something :P

jimthompson5910 (jim_thompson5910):

oh interesting, never seen or heard of that before

OpenStudy (anonymous):

actaully some question written in australia are written 8r9 which means 8/9

jimthompson5910 (jim_thompson5910):

i gotcha, that makes more sense now

jimthompson5910 (jim_thompson5910):

never ever seen that notation before